# SAINT: Attention-Based Modeling of Sub-Action Dependencies in Multi-Action Policies

### v1

@@ 5 Experimental Evaluation
In our experiments, we address three primary research questions: (1) How effective is SAINT in environments with primarily state-independent sub-action dependencies? (2) How effective is SAINT in environments with primarily state-dependent sub-action dependencies? (3) How effective is SAINT in environments with weak sub-action dependencies but complex dynamics? We address these questions in Sections 5.15.2, and 5.3, respectively. In Section 5.4, we present ablation studies that compare FiLM-based conditioning to alternative state injection strategies and measure the computational cost of SAINT’s architecture.
We compare SAINT to four baselines: (1) a pure factorization approach with independent sub-action policies, (2) an autoregressive model [47], (3) Wol-DDPG [8], and (4) a standard RL algorithm selected based on best empirical performance in each environment (PPO in Sections 5.1 and 5.3; A2C in Section 5.2). All reported results are averaged over five random seeds.
@@ 5.1 State-Independent Sub-Action Dependencies
To evaluate SAINT in environments where sub-action dependencies are primarily state-independent, we use the CityFlow traffic control benchmark [46], where each action corresponds to simultaneous phase decisions across multiple intersections. While coordination is necessary to achieve global traffic efficiency, the structure of these dependencies remains largely unchanged across states.
@@
We adopt PPO as the standard RL algorithm in this setting. SAINT, the pure factorized baseline, and the AR baseline, use the same PPO implementation and hyperparameters, isolating the impact of architectural differences. Wol-DDPG performs poorly in this environment and is excluded from the main learning curves; full training curves, including Wol-DDPG, are included in Appendix A.2. Wol-DDPG’s ineffectiveness aligns with prior observations that it is ill-suited to environments with unordered sub-actions [5].
@@
Figure 2: Learning curves show that SAINT outperforms all baselines in both learning speed and final reward, across both the CityFlow-Linear and CityFlow-Irregular environments. In CityFlow-Irregular, SAINT’s relative advantage widens as coordination demands increase.
As can be seen in Figure 2, SAINT learns faster and achieves higher final performance than all baselines in both environments. The gains are especially pronounced in CityFlow-Irregular, where the increased coordination complexity makes the limitations of factorized and sequential approaches more apparent. Notably, PPO, which treats the entire joint action space monolithically, fails to learn a policy approaching the performance of SAINT in either setting.
@@ 5.2 State-Dependent Sub-Action Dependencies
Next, we evaluate SAINT in environments where sub-action dependencies are strongly state-dependent. In these domains, the relationships among sub-actions vary significantly with the environment state, requiring policies to model context-sensitive joint decisions. To study this setting, we use the Combinatorial Navigation Environment (CoNE) [20], a configurable high-dimensional control domain designed to evaluate policy architectures under large, discrete, and structured action spaces.
The agent begins at a fixed origin $s_{0}$ and must reach a predefined goal $g$. At each timestep, it selects a joint action by activating multiple discrete sub-actions, each corresponding to movement along a distinct dimension of the environment. These sub-actions are executed in parallel to produce a single composite transition. The agent receives a reward $r=-\rho(s,g)$ based on the Euclidean distance to the goal. Episodes terminate either upon reaching the goal (reward $+10$) or entering a terminal failure state (pit), which incurs a penalty of $r=-10\cdot\rho(s_{0},g)$ to discourage reward hacking through early failure.
In CoNE, both the action and state spaces grow exponentially with dimensionality: the number of joint actions scales as $|\mathcal{A}|=2^{2D}$, and the number of states as $|\mathcal{S}|=M^{D}$ in a $D$-dimensional environment with $M$ positions per axis. In our largest setting, the environment comprises over 200 million states and nearly 17 million joint actions per state. Beyond scale, CoNE introduces strong sub-action dependencies: some combinations enable efficient movement, others cancel out, and some lead to catastrophic failure. Crucially, these sub-action interactions are highly state-dependent; a combination that is optimal in one state may lead to a pit in another, making effective decision-making highly sensitive to global context. CoNE is highly configurable, allowing us to systematically vary the number of dimensions and pit density to assess the impact of increasing action space size and sub-action dependence.
@@
The results in Table 1 show that SAINT maintains strong performance as action dimensionality increases, significantly outperforming all baselines at every scale. While the factorized and autoregressive baselines achieve modest performance in lower dimensions, their performance degrades or plateaus as the number of sub-actions grows. SAINT’s relative advantage increases in the largest settings, where it maintains low variance and stable performance. Wol-DDPG and A2C perform poorly throughout, likely due to their inability to scale to large, unordered action spaces. Note that A2C is omitted at the highest dimensionality due to the computational intractability of modeling the full joint action space with a flat categorical policy. Full learning curves are provided in Appendix B.2.
@@ 5.3 Weak Sub-Action Dependencies with Complex Dynamics
To evaluate SAINT in environments where sub-action dependencies are relatively weak but the underlying dynamics are complex, we consider discretized variants of the HalfCheetah, Hopper, and Walker2D MuJoCo locomotion tasks [39]. In these environments, each continuous joint control signal is discretized into 11 bins, yielding large, structured action spaces while retaining the rich temporal and physical dynamics of the original tasks. Although the discretized action spaces are combinatorially large, prior work [3] suggests that the dependencies among sub-actions are relatively weak in these domains. This setting thus serves as a testbed for assessing whether SAINT’s architectural benefits persist even when strong coordination between sub-actions is not required.
Given the limitations of PPO and Wol-DDPG identified in Sections 5.1 and 5.2, we limit our comparison to the pure factorized and autoregressive baselines, using identical PPO implementations to ensure a controlled evaluation.
@@
As shown in Figure 3, SAINT performs comparably to the factorized and autoregressive baselines in HalfCheetah. In Hopper and Walker2D, SAINT achieves faster learning and higher final returns, indicating that modeling sub-action structure can provide advantages even in settings with relatively weak dependencies.
5.4 Ablations
To assess the effect of SAINT’s architecture, we conduct targeted ablation studies. First, we compare SAINT’s pre-attention FiLM-based state conditioning to several alternative strategies for injecting state information. Second, we analyze the computational cost of SAINT relative to a pure factorization baseline, quantifying the trade-off between representational capacity and efficiency. These evaluations are performed in the CityFlow-Irregular environment and in the 10-dimensional CoNE setting with pits occupying 25% of interior states.
@@ 5.4 Ablations
We compare SAINT’s FiLM-based state conditioning to several alternative mechanisms for injecting global state information. As shown in Appendix C, FiLM yields higher final performance and more stable training in CityFlow, and performs at least as well as other methods in CoNE. These results suggest that although several conditioning mechanisms are effective, FiLM provides a performance advantage and may promote more stable training dynamics.
Attention Cost
To isolate the computational overhead of modeling sub-action dependencies via self-attention, we compare SAINT’s runtime to that of the pure factorization baseline. We also evaluate a variant of SAINT, called SAINT-IP, that replaces standard self-attention with an inducing point mechanism [21], which approximates full attention using a fixed set of learned summary vectors. This technique reduces the quadratic cost of attention by attending first from the inducing points to the inputs, and then from the inputs back to the summaries. All experiments were conducted on a single NVIDIA A40 GPU using Python 3.9 and PyTorch 2.6. We report wall-clock time per training episode in seconds, averaged over 5 runs.
@@
[…]
@@
[…]
Table 3: Total wall-clock training time and time to reach the factored baseline’s final performance, both measured in seconds. "Total training time" reflects the full duration of training, while "time to factored performance" measures how quickly each method reaches that baseline.
As shown in Table 3, SAINT incurs higher total training time than the pure factorization baseline, reflecting the additional cost of modeling sub-action dependencies via self-attention. SAINT-IP further increases this overhead due to the constant-factor cost of the Induced Set Attention Block (ISAB), which performs two full attention passes per layer. Standard self-attention (used in SAINT), by contrast, requires only one. Although the number of sub-actions in our environments is nontrivial, it remains small relative to domains like large language modeling or 3-D vision, for which attention-based inducing points were originally designed [21]. Consequently, the quadratic cost of full self-attention is not prohibitive, and ISAB’s asymptotic advantage does not yield runtime benefits in practice.
Although self-attention introduces per-step overhead, both SAINT and SAINT-IP reach the factored baseline’s asymptotic performance significantly faster. In CityFlow, they do so in less than one-third of the total training time; in CoNE, SAINT reaches this level nearly 30% faster. These results underscore SAINT’s practical efficiency: while modeling dependencies incurs additional cost, it leads to faster acquisition of strong policies. As shown in Appendix D, SAINT-IP exhibits some degradation in final performance relative to SAINT in CityFlow, but still outperforms the pure factorization baseline in both environments.
@@ 6 Discussion and Conclusion
We introduce SAINT, a Transformer-based policy architecture for combinatorial action spaces with interdependent sub-actions. Unlike existing approaches that assume independence or impose fixed autoregressive orderings, SAINT treats actions as unordered sets and models their structure using permutation-invariant self-attention. A FiLM-based global conditioning mechanism allows SAINT to incorporate state context efficiently, enabling it to capture rich, state-dependent sub-action interactions without sacrificing tractability.
While SAINT demonstrates strong performance and scalability in complex combinatorial action spaces, its limitations present opportunities for future work. First, the self-attention mechanism, while effective for modeling sub-action dependencies, introduces higher per-step computational cost than purely factorized baselines. Although Section 5.4 shows this is often offset by improved sample efficiency and faster convergence, lighter-weight attention variants (e.g., sparse attention or optimized ISAB for large $A$) could benefit resource-constrained settings or extremely high-dimensional action spaces. Second, our formulation assumes a fully observable MDP. While FiLM-based state conditioning enables adaptation to different contexts, handling significant partial observability may require incorporating memory modules or recurrent structures prior to the FiLM layer. Third, SAINT relies on the FiLM network to inject state information into sub-action embeddings. Although ablations show this to be effective, performance may be sensitive to the complexity of state-action relationships and the capacity of the conditioning network. Finally, while SAINT is designed as a general-purpose architecture and performs well across domains (e.g., traffic, navigation, locomotion), entirely novel tasks with different sub-action structures may require tuning. Future work could explore more robust generalization or meta-learning approaches for rapid adaptation to new combinatorial tasks.
Despite these limitations, extensive experiments show that SAINT consistently outperforms strong baselines in both learning speed and final performance, scaling to action spaces with nearly 17 million joint actions. Ablations confirm the effectiveness of our state injection strategy, and runtime analysis shows that the cost of modeling sub-action dependencies is offset by faster attainment of performant policies.
@@ A.1 Environmental Setup
In both CityFlow-Linear (Figure 4(a)) and CityFlow-Irregular (Figure 4(b)), the state is represented as a flat integer vector, in which each value indicates the number of waiting vehicles on an incoming lane and its paired outgoing lane. The reward at each step is the negative of the average "pressure" across intersections, where an intersection’s pressure is defined as the absolute difference between its total incoming and outgoing vehicle counts. Pressure is a standard metric in traffic signal control literature, commonly used to quantify imbalance in intersection flow [23, 42].

### v2

@@ 5 Experimental Evaluation
Our experiments evaluate the efficacy of different action representations for modeling complex sub-action interactions across three regimes: (1) primarily state-independent interactions, (2) state-dependent interactions, and (3) weak interactions with complex dynamics. Results are presented in Sections 5.15.2, and 5.3, respectively. Section 5.4 evaluates SAINT in the offline RL setting. Section 5.5 analyzes key architectural choices, quantifying the trade-off between representational power and computational cost, and robustness to Transformer structural parameters.
We compare SAINT to four baselines reflecting the standard representational assumptions for combinatorial action spaces: (1) a factorized policy (Tavakoli et al., 2018), assuming fully independent sub-actions; (2) an autoregressive model (Zhang et al., 2018), imposing a fixed sequential order; (3) Wol-DDPG (Dulac-Arnold et al., 2015), using a continuous embedding; and (4) a flat RL algorithm, which learns a monolithic representation of the full action space without exploiting its combinatorial structure. Results are averaged over five random seeds.
These four baselines instantiate the dominant structural assumptions used to scale RL to large combinatorial action spaces. The flat policy corresponds to a monolithic model that ignores compositional structure and treats each joint action as an atomic symbol. The factorized policy enforces independent per-dimension decisions, preventing it from modeling necessary coordination between sub-actions. The autoregressive model imposes a fixed sequential ordering over sub-actions, which can be misaligned with the true, permutation-invariant dependency structure. Wol-DDPG embeds each joint action as a single continuous vector, collapsing the internal structure needed to capture interactions among sub-actions. Our experiments in Sections 5.1-5.4 are designed to test whether these structural assumptions remain sufficient when sub-action indexing is arbitrary or only weakly meaningful, or whether a set-based alternative such as SAINT is required.
Because a subset of large joint-action problems are naturally formulated as multi-agent RL, we additionally evaluated a centralized training, decentralized execution (CTDE) baseline — the predominant paradigm in multi-agent RL (Iqbal and Sha, 2019; Rashid et al., 2020; Yu et al., 2022). As reported in Appendix D, CTDE did not outperform the fully factorized policy and was less stable during training. This suggests that performance is primarily governed by execution-time policy expressivity (i.e., the ability to model coordinated joint-action distributions) rather than by adding global information to the critic.
@@ 5.1 State-Independent Sub-Action Dependencies
To evaluate SAINT in environments where sub-action dependencies are primarily state-independent, we use the CityFlow traffic control benchmark (Zhang et al., 2019), where each action corresponds to simultaneous phase decisions across multiple intersections. While coordination is necessary to achieve global traffic efficiency, the structure of these dependencies remains largely unchanged across states.
@@
We adopt PPO as the standard RL algorithm in this setting. SAINT, the pure factorized baseline, and the AR baseline, use the same PPO implementation and hyperparameters, isolating the impact of architectural differences. Wol-DDPG performs poorly in this environment and is excluded from the main learning curves; full training curves, including Wol-DDPG, are included in Appendix A.2. Wol-DDPG’s ineffectiveness aligns with prior observations that it is ill-suited to environments with unordered sub-actions (Chen et al., 2023).
The results in Figure 2, demonstrate the advantage of SAINT’s set-based action representation. In both environments, SAINT learns faster and achieves higher final performance. The factorized approach is structurally incapable of modeling coordination between intersections and thus performs poorly. The autoregressive model is restricted by its fixed sequential prior, which is misaligned with the unordered nature of traffic-signal control. The flat PPO baseline, forced to learn a monolithic representation without exploiting combinatorial structure, fails to learn an effective policy.
@@
Figure 2: Learning curves show that SAINT outperforms all baselines in both learning speed and final reward, across both the CityFlow-Linear and CityFlow-Irregular environments.
@@ 5.2 State-Dependent Sub-Action Dependencies
Next, we evaluate SAINT in environments where sub-action dependencies are strongly state-dependent. In these domains, the relationships among sub-actions vary significantly with the environment state, requiring policies to model context-sensitive joint decisions. To study this setting, we use the Combinatorial Navigation Environment (CoNE) (Landers et al., 2024), a configurable high-dimensional control domain designed to evaluate policy architectures under large, discrete, and structured action spaces.
The agent begins at a fixed origin $s_{0}$ and must reach a predefined goal $g$. At each timestep, it selects a joint action by activating multiple discrete sub-actions, each corresponding to movement along a distinct dimension of the environment. These sub-actions are executed in parallel to produce a single composite transition. The agent receives a reward $r=-\rho(s,g)$ at each step based on the Euclidean distance to the goal. Episodes terminate either upon reaching the goal (reward +10+10) or entering a terminal failure state (pit), which incurs a penalty of $r=-10\cdot\rho(s_{0},g)$ to discourage reward hacking through early failure.
In CoNE, both the action and state spaces grow exponentially with dimensionality: the number of joint actions scales as $|\mathcal{A}|=2^{2D}$, and the number of states as $|\mathcal{S}|=M^{D}$ in a $D$-dimensional environment with $M$ positions per axis. In our largest setting, the environment comprises over 200 million states and nearly 17 million joint actions per state. Beyond scale, CoNE introduces strong sub-action dependencies — some combinations enable efficient movement, others cancel out, and some lead to catastrophic failure. Crucially, these sub-action interactions are highly state-dependent; a combination that is optimal in one state may lead to a pit in another, making effective decision-making highly sensitive to global context. CoNE is highly configurable, allowing us to systematically vary the number of dimensions and pit density to assess the impact of increasing action space size and sub-action dependence.
@@
The results in Table 1 show that SAINT maintains strong performance as action dimensionality increases, significantly outperforming all baselines at every scale. While the factorized and autoregressive baselines achieve modest performance in lower dimensions, their performance degrades or plateaus as the number of sub-actions grows. This suggests their fixed representational priors — assuming either complete independence or a single fixed order — are insufficient to capture the complex interactions that emerge at scale. SAINT’s relative advantage increases in the largest settings, where it maintains low variance and stable performance. Wol-DDPG and A2C perform poorly throughout, highlighting their inability to form a tractable and meaningful representation of large, unordered action spaces. Note that A2C is omitted at the highest dimensionality due to the computational intractability of modeling the full joint action space with a flat categorical policy. Full learning curves are provided in Appendix B.2.
@@ 5.3 Weak Sub-Action Dependencies with Complex Dynamics
To evaluate SAINT in environments where sub-action dependencies are relatively weak but the underlying dynamics are complex, we consider discretized variants of the HalfCheetah, Hopper, and Walker2D MuJoCo locomotion tasks (Towers et al., 2024). In these environments, each continuous joint control signal is discretized into 11 bins, yielding large, structured action spaces while retaining the rich temporal and physical dynamics of the original tasks. Although the discretized action spaces are combinatorially large, prior work (Beeson et al., 2024) suggests that the dependencies among sub-actions are relatively weak in these domains. This setting thus provides a useful test of SAINT’s generality, elucidating whether the architectural overhead of self-attention remains beneficial when sub-action dependencies are weak or whether simpler factorized policies suffice.
Given the limitations of PPO and Wol-DDPG identified in Sections 5.1 and 5.2, we restrict our comparison to the pure factorized and autoregressive baselines, using identical PPO implementations to ensure a controlled evaluation.
@@
As shown in Figure 3, SAINT matches baseline performance in HalfCheetah and achieves faster learning and higher returns in Hopper and Walker2D. This demonstrates that even when sub-action dependencies are weak, learning a set-based representation of the action space gives an advantage over the rigid assumptions of factorization or a fixed autoregressive order.
5.4 Offline RL
Finally, we evaluate whether SAINT can be used effectively as a policy architecture in offline RL. Specifically, we use the medium-expert datasets from the discretized DM Control tasks cheetah run, finger spin, humanoid stand, quadruped walk, and dog trot introduced by Beeson et al. (2024). Across these environments, the number of sub-actions ranges from six to 39.
In this section, we report results using Batch-Constrained Q-learning (BCQ) (Fujimoto et al., 2019) as the offline RL objective. Additional offline evaluations with AWAC and IQL are provided in Appendix E. As in Section 5.3, we limit our comparison to the pure factorized and autoregressive baselines.
Task SAINT Factored AR
[…]
Table 3: Mean $\pm$ std performance on offline DM Control tasks with BCQ variants.
As shown in Table 3, SAINT-BCQ achieves the strongest performance across all domains, demonstrating that explicitly modeling sub-action interactions improves policy quality even when learning is restricted to logged trajectories.
5.5 Analyses and Ablations
To evaluate SAINT’s design choices and robustness, we conduct three analyses. First, we compare FiLM-based state conditioning to alternative mechanisms. Second, we assess the trade-off between representational power and computational cost, quantifying how modeling sub-action interactions affects sample efficiency. These evaluations are performed in the CityFlow-Irregular environment and in the 10-dimensional CoNE setting with pits occupying 25% of interior states. Finally, we test SAINT’s robustness to architectural hyperparameters using the CityFlow-Irregular environment.
@@ 5.5 Analyses and Ablations
We compare SAINT’s FiLM-based state conditioning to four alternatives: (1) applying cross-attention to the state before self-attention, (2) applying cross-attention to the state after self-attention, (3) interleaving cross-attention and self-attention layers, and (4) appending the state as an additional token within the sub-action self-attention block. As shown in Appendix C, FiLM achieves higher final performance and more stable training in CityFlow, and performs at least as well as the alternatives in CoNE. These results indicate that while multiple conditioning mechanisms are effective, FiLM provides a consistent performance advantage and stabilizes training.
Representational Power vs. Sample Efficiency
To isolate the computational overhead of modeling sub-action dependencies via self-attention, we compare SAINT’s runtime to that of the pure factorization baseline. We also evaluate a variant of SAINT, called SAINT-IP, that replaces standard self-attention with an inducing point mechanism (Lee et al., 2019), which approximates full attention using a fixed set of learned summary vectors. This technique reduces the quadratic cost of attention by attending first from the inducing points to the inputs, and then from the inputs back to the summaries. All experiments were conducted on a single NVIDIA A40 GPU using Python 3.9 and PyTorch 2.6. We report wall-clock time per training episode in seconds, averaged over 5 runs.
As shown in Table 4, SAINT requires more training time than the pure factorization baseline, reflecting the added cost of modeling sub-action dependencies with self-attention. SAINT-IP incurs further overhead from the Induced Set Attention Block (ISAB), which performs two attention passes per layer, compared to one in standard self-attention. While the number of sub-actions in our environments is nontrivial, it remains small relative to domains such as large language modeling or 3-D vision, for which inducing points were introduced (Lee et al., 2019). Consequently, the quadratic cost of full self-attention is not prohibitive, and ISAB’s asymptotic advantage does not yield runtime benefits.
However, in practical settings efficiency is better measured by wall-clock time to reach a target return. The “time to factored performance” metric shows that SAINT and SAINT-IP reach the factorized baseline’s asymptotic return in less than one-third of the time in CityFlow and about 30% faster in CoNE. This indicates that the added computation of explicit action representations is outweighed by faster learning — by modeling dependency structure, SAINT achieves stronger policies in less time.
@@
   Total training time $5204.2$ $5412.4$ 3936.2
   Time to factored performance 1088.53 1087.49 $3936.2$
@@
   Total training time $966.5$ $980.9$ 535.9
   Time to factored performance 395.76 $504.97$ $535.9$
Table 4: Total wall-clock training time and time to reach the factored baseline’s final performance, both measured in seconds. “Total training time” reflects the full duration of training, while “time to factored performance” measures how quickly each method reaches that baseline.
Robustness to Architectural Hyperparameters
We evaluated SAINT’s sensitivity to architectural hyperparameters by sweeping the number of self-attention blocks $\{1,3,5\}$ and attention heads $\{1,2,4,8\}$, for a total of twelve configurations. As shown in Appendix G, performance varied within a narrow range — the best setting (3 blocks $\times$ 1 head) outperformed the weakest (1 block $\times$ 8 heads) by only $\sim 7\%$. Even the weakest configuration exceeded all baselines, underscoring SAINT’s robustness to attention depth and head count.
@@ 6 Discussion and Conclusion
We introduce SAINT, a policy architecture that treats learning in large discrete action spaces as a representation learning problem. Instead of assuming conditional independence or a fixed ordering, SAINT learns a state-conditioned, permutation-equivariant set representation of the combinatorial action space. Self-attention models the interactions within this set, yielding expressive and tractable policies for combinatorial domains.
While SAINT achieves strong performance, several limitations remain. First, while section 5.5 shows that the cost of self-attention is often offset by improved sample efficiency, lighter-weight attention variants such as sparse attention (Child et al., 2019) could benefit resource-constrained settings. Second, our analyses also validate FiLM as an effective state-conditioning mechanism, yet performance in new domains may depend on the capacity of this network, motivating exploration of more expressive state-injection methods such as those proposed in multi-agent RL (Iqbal and Sha, 2019). Third, SAINT is designed for domains wherein a joint action is naturally represented as a set of parallel sub-actions for which indexing is arbitrary or weakly structured. In settings with highly structured and known priors, a fully permutation-equivariant prior may not be the most effective representation. In such settings hybrid architectures that combine structured embeddings with partial equivariance are a promising direction for future work. Finally, although our experiments assume a fixed set of sub-actions, many real-world domains, such as road closures or reconfigurable network topologies, involve dynamically changing action sets. Because SAINT represents actions as an unordered set, it naturally supports such variability via masking. Systematically evaluating this capability is an important research direction.
In this work we show that learning explicit representations of sub-action interactions is an effective and practical approach for control in combinatorial action spaces. SAINT represents actions as unordered sets and applies self-attention to capture sub-action dependencies, yielding expressive yet tractable policies. This modeling delivers substantial gains in sample efficiency, accelerating convergence to high-performing policies. In environments with up to 17 million joint actions, SAINT consistently outperforms baselines that assume independence, impose ordering, or learn flat policies, demonstrating the effectiveness of modeling sub-action interactions for scalable combinatorial control.
@@ A.1 Environmental Setup
In both CityFlow-Linear (Figure 4(a)) and CityFlow-Irregular (Figure 4(b)), the state is represented as a flat integer vector, in which each value indicates the number of waiting vehicles on an incoming lane and its paired outgoing lane. The reward at each step is the negative of the average “pressure” across intersections, where an intersection’s pressure is defined as the absolute difference between its total incoming and outgoing vehicle counts. Pressure is a standard metric in traffic signal control literature, commonly used to quantify imbalance in intersection flow Lioris et al. (2016); Varaiya (2013).